OpenStudy (smurfy14):

Find the real numbers x and y (x+6i)=(3-i)+(4-2yi) so far i have x+6i=7... and idk what to do next

OpenStudy (anonymous):

3.141592654 is the answer. But plug it into the migrator then find the binomial coefficient first.

jimthompson5910 (jim_thompson5910):

x+6i=(3-i)+(4-2yi) x+6i=3-i+4-2yi x+6i=7+(-1-2y)i x+6i=7-(1+2y)i Equate corresponding real and imaginary parts. So x = 7 and 6i = -(1+2y)i, which means that 6 = -(1+2y). Solve for y to get y = -7/2